OpenStudy (anonymous):

Solve for x in the equation: 16x^3 + 4x^2 - 4x -1 (all solutions are rational and between +- 1

OpenStudy (anonymous):

Factor by grouping, then use the zero product property.

OpenStudy (radar):

You yet to have an equation where is the = sign?

OpenStudy (anonymous):

Good catch, radar, I assumed equal to zero....

OpenStudy (radar):

I bet you're right @AnimalAin

OpenStudy (anonymous):

...and we all know where assumptions get us...

OpenStudy (radar):

Yes, I did not want to break it down that far lol

OpenStudy (radar):

If it is equal to zero, one of the roots might be 1/2

OpenStudy (radar):

In fact it is by inspection.

OpenStudy (radar):

If it were equal to zero that is.

OpenStudy (radar):

Continuation of this assumption and dividing the expression by (x-1/2) gives us a quadratic of\[16x ^{2}+12x+2\]

OpenStudy (anonymous):

On the summer work assignment that my teacher created it doesn't have an equal sign, yet does say equation,so do I assume or not?

OpenStudy (anonymous):

I would think so. In that case,\[16x^3 + 4x^2 - 4x -1=0 \implies 4x^2(4x+1)-(4x+1)=0\]\[\implies 16(x+\frac{1}{2})(x-\frac{1}{2})(x+\frac{1}{4})=0 \implies x \in \left\{ -\frac{1}{2},-\frac{1}{4},\frac{1}{2} \right\}\]

OpenStudy (radar):

well it won't let me use the editor. 16x^2 +12x +2 =0 factor that and get more roots. Your teacher has put you in a position that you probably have to assume the expression is equal to 0. Chastise her for me when you see her.

OpenStudy (anonymous):

From Mathematica:\[16 x^3+4 x^2-4 x-1=(2 x-1) (2 x+1) (4 x+1) \]

OpenStudy (anonymous):

Lills, you understand this?

OpenStudy (anonymous):

I see where your answer is coming from, but I'm not sure if given another problem I would be able to do it alone

OpenStudy (anonymous):

Look for the patterns. In this case, we have a cubic with coefficients that allow factoring by grouping. After that, we have one factor that is a difference of squares. When the whole kebab is factored, you just use the zero product property.

OpenStudy (anonymous):

What I am not completely following is how you got from the original equation to the next step, I understand the zero part

OpenStudy (anonymous):

Note that the ratio between the first and second coefficients is the same as the ratio between the third and fourth coefficients. This indicates we can factor by grouping. A little more on this here http://coolmath.com/algebra/04-factoring/07-factor-by-grouping-01.htm or about halfway down the page here http://tutorial.math.lamar.edu/Classes/Alg/Factoring.aspx

OpenStudy (anonymous):

Thanks you so much